# Measuring Social Norms of Large Language Models

### v1

@@ Appendix D Summary of Skills
Description: "Vociferous" and "vocal" are related English words that come from Latin. What does **vociferous** mean?
Choices: [(a) loud and forceful, (b) wise and thoughtful, (c) strong and fit]
Answer index: 0
Subject: Language Arts
Skill: analyze-passages-from-the-lightning-thief-part-2
Description: Read the following passage from "The Lightning Thief". In this excerpt, Percy Jackson, the narrator, is describing a dream.
In my dreams, I stood in a dark cavern before a gaping pit. Gray mist creatures churned all around me, whispering rags of smoke that I somehow knew were the spirits of the dead.
From Rick Riordan, The Lightning Thief. Copyright 2005 by Rick Riordan
What is most likely true about the spirits in Percy’s dream?
Choices: [(a) The spirits will cooperate with the ancient voice., (b) The spirits are trying to protect Percy., (c) The spirits want to lead Percy into the pit.]
Answer index: 1
Subject: Language Arts
Skill: analyze-passages-from-harriet-tubman-conductor-on-the-underground-railroad-part-1
Description: Review the passage.
@@
Harriet thought that Christmas was the very best time of all. By tradition there was no work. The holiday for the field slaves lasted as long as the Yule log burned in the fireplace at the Big House. So the people in the quarter spent days preparing the log. They chose a big one, so big that the strongest field hands bent their backs under its weight. They soaked it in water, so that it would burn slowly and for a long time.
Adapted from Ann Petry, Harriet Tubman: Conductor on the Underground Railroad. Copyright 1955 by Ann Petry
Which item does the author use to symbolize comfort and security?
Choices: [(a) the cabin, (b) the overseer’s horn, (c) the blankets]
Answer index: 0
Subject: Language Arts
Skill: analyze-passages-from-roll-of-thunder-hear-my-cry-part-1
Description: Review the passage.
"Hey, Cassie," said Mary Lou Wellever, the principal’s daughter, as she flounced by in a new yellow dress.
From Mildred D. Taylor, Roll of Thunder, Hear My Cry. Copyright 1976 by Mildred D. Taylor
How do students feel about wearing shoes on the first day of school?
Choices: [(a) They are embarrassed by their shoes., (b) They are eager to wear their shoes., (c) They are uncomfortable in their shoes.]
Answer index: 2
Subject: Language Arts
Skill: analyze-passages-from-roll-of-thunder-hear-my-cry-part-2
Description: Review the passage.
@@
The revival ran for seven days and it was an occasion everyone looked forward to, for it was more than just church services; it was the year’s only planned social event, disrupting the humdrum of everyday country life. Teenagers courted openly, adults met with relatives and friends they had not seen since the previous year’s "big meeting," and children ran almost free.
From Mildred D. Taylor, Roll of Thunder, Hear My Cry. Copyright 1976 by Mildred D. Taylor
Look at the text in **bold**. What does this figurative language show about the revival?
Choices: [(a) It shows how long and drawn out the service is., (b) It shows how energetic the townspeople are., (c) It shows how overcrowded the event is.]
Answer index: 2
Subject: Language Arts
Skill: draw-inferences-from-literary-texts
Description: Read the text.
From his seat at a table in the local library, Ruben turned to the side and saw Nora on a couch, deeply engrossed in a fantasy book. Nora barely seemed to blink as she read, and Ruben thought he had never seen anyone so absorbed in a story.
Suddenly, Ruben felt a rumbling beneath his feet. The model of the solar system suspended from the ceiling began swinging back and forth, as books started falling off the nearby shelves. There was a commotion as library patrons hustled out of the way.
Nora stood up and in a clear, loud voice, shouted, "Everybody! This is an earthquake. Get under the tables, now!"
Everyone looked at Nora in shock, frozen in place by a newfound fear.
"Do it!" Nora commanded, as the floor shifted. "Do it right now!"
Ruben and the other patrons followed Nora’s instructions, scrambling under the library tables.
"Now, put your hands over your head and neck, face away from the windows, and stay still," Nora instructed.
As the room shook, some people squealed in fear, while others made soft whimpering noises.
From underneath her own table, Nora calmly reassured them. "Don’t worry—this one doesn’t seem too bad, and hopefully it’ll be over in just a few seconds."
What is most likely true about Nora?
Choices: [(a) She tends to overreact in stressful situations., (b) She has prior experience with earthquakes., (c) She works at the library and knows the patrons well.]
Answer index: 1
Subject: Language Arts
Skill: analyze-passages-from-the-lightning-thief-part-1
Description: Reread the following passage from "The Lightning Thief". In this excerpt, Percy is talking with his friend Annabeth after fighting a tough battle on their quest.
I was pretty much in shock myself. The explosion of bus windows still rang in my ears. But Annabeth kept pulling us along, saying: "Come on! The farther away we get, the better."
From Rick Riordan, The Lightning Thief. Copyright 2005 by Rick Riordan
Which of the following is a reason Annabeth does not want Percy to die?
Choices: [(a) If Percy dies, Annabeth would probably have to live with her dad., (b) If Percy dies, Annabeth would probably have to finish the quest on her own., (c) If Percy dies, Annabeth would probably have to go back to the camp.]
Answer index: 2
Subject: Language Arts
Skill: evaluate-newspaper-headlines-for-bias
Description: Which headline best avoids biased language?
@@
Choices: [(a) Second Grader’s Peach Pie Wins First Place in Baking Contest, (b) Inexplicably, Judges Pick Child’s Peach Pie Despite Tastier Entries, (c) Well-Earned Prize Goes to Young Student’s Mouthwatering Pie]
Answer index: 0
@@
Subject: Language Arts
Skill: analyze-passages-from-harriet-tubman-conductor-on-the-underground-railroad-part-2
Description: Review the passage.
@@
In the fall of 1854, Harriet Tubman began to feel uneasy about three of her brothers. Benjamin, John and William Henry were still in Maryland, working on plantations where they had been hired out. She kept having dreams about them, vivid dreams in which she saw them sold and sent away in a chain gang. She decided to tell them that she was coming to Maryland that fall, so that they would be ready to go North with her.
Adapted from Ann Petry, Harriet Tubman: Conductor on the Underground Railroad. Copyright 1955 by Ann Petry
What is the meaning of the word **eluded** as used in the passage?
Choices: [(a) escaped the understanding of, (b) caused excitement about, (c) raised awareness for]
Answer index: 0
Subject: Language Arts
Skill: classify-the-figure-of-speech-euphemism-hyperbole-oxymoron-paradox
Description: Which figure of speech is used in this text?
Lauren asked her mother if she could adopt a cat, and her mother replied, "It’s a **definite maybe**," so Lauren didn’t want to get her hopes up.
Choices: [(a) euphemism, (b) oxymoron]
Answer index: 1
Subject: Language Arts
Skill: classify-the-figure-of-speech-anaphora-antithesis-apostrophe-assonance-chiasmus-understatement
Description: Which figure of speech is used in this text?
Soft language **issued** from their **spittleless** **lips** as they **swished** **in** slow circles round and round the field.
—James Joyce, "A Portrait of the Artist as a Young Man"
Choices: [(a) chiasmus, (b) assonance]
Answer index: 1
Subject: Language Arts
Skill: interpret-the-figure-of-speech
Description: What does the **idiom** in this text suggest?
"Speak of the devil!" Trudy declared when Bert strolled into the room.
Choices: [(a) Trudy had just been speaking about Bert., (b) Trudy didn’t trust Bert.]
Answer index: 0
Subject: Language Arts
Skill: prefixes
Description: Why would a company require an employee to sign a **nondisclosure** agreement?
@@
Choices: [(a) to keep company secrets private, (b) to prevent the employee from working for a competitor]
Answer index: 0
Subject: Language Arts
Skill: suffixes
Description: If an objection is **meritless**, what should you do?
Choices: [(a) ignore or dismiss it, (b) take it seriously]
Answer index: 0
Subject: Language Arts
Skill: domain-specific-vocabulary-in-context-science-and-technical-subjects
Description: Read the passage and then answer the question.
@@
"Rafflesia arnoldii" produces the largest flowers of any plant known to man. It also lacks leaves, roots, a stem, and functioning chloroplasts. What exactly is this bizarre organism, and how does it survive without so many seemingly crucial parts? **Endemic** to the rainforests of Southeast Asia, """R. arnoldii""" has developed a number of adaptations that enable it to thrive in its tropical home. Key to its survival is its parasitic relationship with a grapelike vine of the genus "Tetrastigma". """R. arnoldii""" spends most of its life almost completely concealed inside the host vine. When it is ready to reproduce, however, it grows a single flower that measures about three feet in diameter, weighs up to fifteen pounds, and emits a strong odor of rotting flesh. The smell of the flower, which attracts carrion flies and other pollinators, has earned """R. arnoldii""" the nickname "corpse flower."
What is the meaning of **endemic** as used in the passage?
Choices: [(a) new, (b) crucial, (c) harmful, (d) native]
Answer index: 3
Subject: Language Arts
Skill: classify-the-figure-of-speech-review
Description: Which figure of speech is used in this text?
His guardians were extremely old-fashioned people and did not realize that **we live in an age when unnecessary things are our only necessities**.
—Oscar Wilde, "The Picture of Dorian Gray"
Choices: [(a) paradox, (b) chiasmus]
Answer index: 0
Subject: Language Arts
Skill: analyze-passages-from-a-long-walk-to-water-part-2
Description: Read the following passage from "A Long Walk to Water". In this excerpt, Salva has just arrived in Rochester, New York, and is meeting his host family.
There they were, smiling and waving in the airport lobby—his new family! Chris, the father; Louise, the mother; and four children. Salva would have siblings, just as he had before. He felt his shoulders relax a little on seeing their eager smiles.
From Linda Sue Park, A Long Walk to Water. Copyright 2010 by Linda Sue Park
How does Salva feel upon meeting his new family?
Choices: [(a) He feels unsure of how to fit into his new family, but he is excited to get to know them., (b) His tension and nerves begin to subside, but he still feels a little overwhelmed., (c) He is thankful for their generosity, but he wants to remain emotionally guarded.]
Answer index: 1
Subject: Language Arts
Skill: analyze-passages-from-a-night-to-remember-part-1
Description: Review the passage.
@@
Up on the bridge, Captain Smith tried to piece the picture together. No one was better equipped to do it. After 38 years’ service with White Star, he was more than just senior captain of the line; he was a bearded patriarch, worshipped by crew and passengers alike. They loved everything about him—especially his wonderful combination of firmness and urbanity. It was strikingly evident in the matter of cigars. "Cigars," says his daughter, "were his pleasure. And one was allowed to be in the room only if one was absolutely still, so that the blue cloud over his head never moved."
From Walter Lord, A Night To Remember. Copyright 1955 by Walter Lord.
Look at the text in **bold**. What is the effect of this sentence on the tone of the passage?
Choices: [(a) It creates a skeptical tone., (b) It creates a somber tone., (c) It creates a frustrated tone.]
Answer index: 1
Subject: Language Arts
Skill: analyze-passages-from-anne-of-green-gables-part-1
Description: Read the following passage from "Anne of Green Gables". In this excerpt, Anne is chatting with her friend Diana at a tea party.
@@
"We had a plum pudding for dinner on Tuesday and there was half the pudding and a pitcherful of sauce left over. Marilla said there was enough for another dinner and told me to set it on the pantry shelf and cover it. I meant to cover it just as much as could be, Diana, but I forgot all about it. I thought of it next morning and ran to the pantry. Diana, fancy if you can my extreme horror at finding a mouse drowned in that pudding sauce! I lifted the mouse out with a spoon and threw it out in the yard. Marilla was out milking and I fully intended to ask her when she came in if I’d give the sauce to the pigs; but when she did come in I was imagining that I was a frost fairy going through the woods turning the trees red and yellow, so I never thought about the pudding sauce again.
Adapted from L. M. Montgomery, Anne of Green Gables
What does this passage suggest about Anne?
Choices: [(a) She does not take life very seriously and wishes that others would relax., (b) She has good intentions, but she can be absentminded sometimes., (c) She is self-confident and does not worry about what other people think.]
Answer index: 1
Subject: Language Arts
Skill: analyze-passages-from-anne-frank-the-diary-of-a-young-girl-part-2
Description: Review the passage.
Think of it, having to sit in such terror for a day and two nights! We thought of nothing, but simply sat there in pitch darkness—in her fear, Mrs. van D. had switched off the lamp. We whispered, and every time we heard a creak, someone said, "Shh, shh."
From Anne Frank, The Diary of a Young Girl: The Definitive Edition. Copyright 1991 by The Anne Frank Fonds. Trans. Susan Massotty
What is the meaning of **engrossed** as it is used in the passage?
Choices: [(a) preoccupied or captivated, (b) puzzled or confused, (c) absentminded or forgetful]
Answer index: 0

### v2

@@ Appendix D Summary of Skills
Subject: Language Arts Skill: use-etymologies-to-determine-the-meanings-of-words Description: "Vociferous" and "vocal" are related English words that come from Latin. What does **vociferous** mean? Choices: [(a) loud and forceful, (b) wise and thoughtful, (c) strong and fit] Answer index: 0
Subject: Language Arts Skill: analyze-passages-from-the-lightning-thief-part-2 Description: Read the following passage from "The Lightning Thief". In this excerpt, Percy Jackson, the narrator, is describing a dream. In my dreams, I stood in a dark cavern before a gaping pit. Gray mist creatures churned all around me, whispering rags of smoke that I somehow knew were the spirits of the dead. From Rick Riordan, The Lightning Thief. Copyright 2005 by Rick Riordan What is most likely true about the spirits in Percy’s dream? Choices: [(a) The spirits will cooperate with the ancient voice., (b) The spirits are trying to protect Percy., (c) The spirits want to lead Percy into the pit.] Answer index: 1
@@
Subject: Language Arts Skill: analyze-passages-from-harriet-tubman-conductor-on-the-underground-railroad-part-1 Description: Review the passage. Harriet thought that Christmas was the very best time of all. By tradition there was no work. The holiday for the field slaves lasted as long as the Yule log burned in the fireplace at the Big House. So the people in the quarter spent days preparing the log. They chose a big one, so big that the strongest field hands bent their backs under its weight. They soaked it in water, so that it would burn slowly and for a long time. Adapted from Ann Petry, Harriet Tubman: Conductor on the Underground Railroad. Copyright 1955 by Ann Petry Which item does the author use to symbolize comfort and security? Choices: [(a) the cabin, (b) the overseer’s horn, (c) the blankets] Answer index: 0
Subject: Language Arts Skill: analyze-passages-from-roll-of-thunder-hear-my-cry-part-1 Description: Review the passage. "Hey, Cassie," said Mary Lou Wellever, the principal’s daughter, as she flounced by in a new yellow dress. From Mildred D. Taylor, Roll of Thunder, Hear My Cry. Copyright 1976 by Mildred D. Taylor How do students feel about wearing shoes on the first day of school? Choices: [(a) They are embarrassed by their shoes., (b) They are eager to wear their shoes., (c) They are uncomfortable in their shoes.] Answer index: 2
@@
Subject: Language Arts Skill: analyze-passages-from-roll-of-thunder-hear-my-cry-part-2 Description: Review the passage. The revival ran for seven days and it was an occasion everyone looked forward to, for it was more than just church services; it was the year’s only planned social event, disrupting the humdrum of everyday country life. Teenagers courted openly, adults met with relatives and friends they had not seen since the previous year’s "big meeting," and children ran almost free. From Mildred D. Taylor, Roll of Thunder, Hear My Cry. Copyright 1976 by Mildred D. Taylor Look at the text in **bold**. What does this figurative language show about the revival? Choices: [(a) It shows how long and drawn out the service is., (b) It shows how energetic the townspeople are., (c) It shows how overcrowded the event is.] Answer index: 2
Subject: Language Arts Skill: draw-inferences-from-literary-texts Description: Read the text. From his seat at a table in the local library, Ruben turned to the side and saw Nora on a couch, deeply engrossed in a fantasy book. Nora barely seemed to blink as she read, and Ruben thought he had never seen anyone so absorbed in a story. Suddenly, Ruben felt a rumbling beneath his feet. The model of the solar system suspended from the ceiling began swinging back and forth, as books started falling off the nearby shelves. There was a commotion as library patrons hustled out of the way. Nora stood up and in a clear, loud voice, shouted, "Everybody! This is an earthquake. Get under the tables, now!" Everyone looked at Nora in shock, frozen in place by a newfound fear. "Do it!" Nora commanded, as the floor shifted. "Do it right now!" Ruben and the other patrons followed Nora’s instructions, scrambling under the library tables. "Now, put your hands over your head and neck, face away from the windows, and stay still," Nora instructed. As the room shook, some people squealed in fear, while others made soft whimpering noises. From underneath her own table, Nora calmly reassured them. "Don’t worry—this one doesn’t seem too bad, and hopefully it’ll be over in just a few seconds." What is most likely true about Nora? Choices: [(a) She tends to overreact in stressful situations., (b) She has prior experience with earthquakes., (c) She works at the library and knows the patrons well.] Answer index: 1
Subject: Language Arts Skill: analyze-passages-from-the-lightning-thief-part-1 Description: Reread the following passage from "The Lightning Thief". In this excerpt, Percy is talking with his friend Annabeth after fighting a tough battle on their quest. I was pretty much in shock myself. The explosion of bus windows still rang in my ears. But Annabeth kept pulling us along, saying: "Come on! The farther away we get, the better." From Rick Riordan, The Lightning Thief. Copyright 2005 by Rick Riordan Which of the following is a reason Annabeth does not want Percy to die? Choices: [(a) If Percy dies, Annabeth would probably have to live with her dad., (b) If Percy dies, Annabeth would probably have to finish the quest on her own., (c) If Percy dies, Annabeth would probably have to go back to the camp.] Answer index: 2
@@
Subject: Language Arts Skill: evaluate-newspaper-headlines-for-bias Description: Which headline best avoids biased language? Choices: [(a) Second Grader’s Peach Pie Wins First Place in Baking Contest, (b) Inexplicably, Judges Pick Child’s Peach Pie Despite Tastier Entries, (c) Well-Earned Prize Goes to Young Student’s Mouthwatering Pie] Answer index: 0
@@
Subject: Language Arts Skill: analyze-passages-from-harriet-tubman-conductor-on-the-underground-railroad-part-2 Description: Review the passage. In the fall of 1854, Harriet Tubman began to feel uneasy about three of her brothers. Benjamin, John and William Henry were still in Maryland, working on plantations where they had been hired out. She kept having dreams about them, vivid dreams in which she saw them sold and sent away in a chain gang. She decided to tell them that she was coming to Maryland that fall, so that they would be ready to go North with her. Adapted from Ann Petry, Harriet Tubman: Conductor on the Underground Railroad. Copyright 1955 by Ann Petry What is the meaning of the word **eluded** as used in the passage? Choices: [(a) escaped the understanding of, (b) caused excitement about, (c) raised awareness for] Answer index: 0
Subject: Language Arts Skill: classify-the-figure-of-speech-euphemism-hyperbole-oxymoron-paradox Description: Which figure of speech is used in this text? Lauren asked her mother if she could adopt a cat, and her mother replied, "It’s a **definite maybe**," so Lauren didn’t want to get her hopes up. Choices: [(a) euphemism, (b) oxymoron] Answer index: 1
Subject: Language Arts Skill: classify-the-figure-of-speech-anaphora-antithesis-apostrophe-assonance-chiasmus-understatement Description: Which figure of speech is used in this text? Soft language **issued** from their **spittleless** **lips** as they **swished** **in** slow circles round and round the field. —James Joyce, "A Portrait of the Artist as a Young Man" Choices: [(a) chiasmus, (b) assonance] Answer index: 1
Subject: Language Arts Skill: interpret-the-figure-of-speech Description: What does the **idiom** in this text suggest? "Speak of the devil!" Trudy declared when Bert strolled into the room. Choices: [(a) Trudy had just been speaking about Bert., (b) Trudy didn’t trust Bert.] Answer index: 0
@@
Subject: Language Arts Skill: prefixes Description: Why would a company require an employee to sign a **nondisclosure** agreement? Choices: [(a) to keep company secrets private, (b) to prevent the employee from working for a competitor] Answer index: 0
Subject: Language Arts Skill: suffixes Description: If an objection is **meritless**, what should you do? Choices: [(a) ignore or dismiss it, (b) take it seriously] Answer index: 0
@@
Subject: Language Arts Skill: domain-specific-vocabulary-in-context-science-and-technical-subjects Description: Read the passage and then answer the question. "Rafflesia arnoldii" produces the largest flowers of any plant known to man. It also lacks leaves, roots, a stem, and functioning chloroplasts. What exactly is this bizarre organism, and how does it survive without so many seemingly crucial parts? **Endemic** to the rainforests of Southeast Asia, """R. arnoldii""" has developed a number of adaptations that enable it to thrive in its tropical home. Key to its survival is its parasitic relationship with a grapelike vine of the genus "Tetrastigma". """R. arnoldii""" spends most of its life almost completely concealed inside the host vine. When it is ready to reproduce, however, it grows a single flower that measures about three feet in diameter, weighs up to fifteen pounds, and emits a strong odor of rotting flesh. The smell of the flower, which attracts carrion flies and other pollinators, has earned """R. arnoldii""" the nickname "corpse flower." What is the meaning of **endemic** as used in the passage? Choices: [(a) new, (b) crucial, (c) harmful, (d) native] Answer index: 3
Subject: Language Arts Skill: classify-the-figure-of-speech-review Description: Which figure of speech is used in this text? His guardians were extremely old-fashioned people and did not realize that **we live in an age when unnecessary things are our only necessities**. —Oscar Wilde, "The Picture of Dorian Gray" Choices: [(a) paradox, (b) chiasmus] Answer index: 0
Subject: Language Arts Skill: analyze-passages-from-a-long-walk-to-water-part-2 Description: Read the following passage from "A Long Walk to Water". In this excerpt, Salva has just arrived in Rochester, New York, and is meeting his host family. There they were, smiling and waving in the airport lobby—his new family! Chris, the father; Louise, the mother; and four children. Salva would have siblings, just as he had before. He felt his shoulders relax a little on seeing their eager smiles. From Linda Sue Park, A Long Walk to Water. Copyright 2010 by Linda Sue Park How does Salva feel upon meeting his new family? Choices: [(a) He feels unsure of how to fit into his new family, but he is excited to get to know them., (b) His tension and nerves begin to subside, but he still feels a little overwhelmed., (c) He is thankful for their generosity, but he wants to remain emotionally guarded.] Answer index: 1
@@
Subject: Language Arts Skill: analyze-passages-from-a-night-to-remember-part-1 Description: Review the passage. Up on the bridge, Captain Smith tried to piece the picture together. No one was better equipped to do it. After 38 years’ service with White Star, he was more than just senior captain of the line; he was a bearded patriarch, worshipped by crew and passengers alike. They loved everything about him—especially his wonderful combination of firmness and urbanity. It was strikingly evident in the matter of cigars. "Cigars," says his daughter, "were his pleasure. And one was allowed to be in the room only if one was absolutely still, so that the blue cloud over his head never moved." From Walter Lord, A Night To Remember. Copyright 1955 by Walter Lord. Look at the text in **bold**. What is the effect of this sentence on the tone of the passage? Choices: [(a) It creates a skeptical tone., (b) It creates a somber tone., (c) It creates a frustrated tone.] Answer index: 1
@@
Subject: Language Arts Skill: analyze-passages-from-anne-of-green-gables-part-1 Description: Read the following passage from "Anne of Green Gables". In this excerpt, Anne is chatting with her friend Diana at a tea party. "We had a plum pudding for dinner on Tuesday and there was half the pudding and a pitcherful of sauce left over. Marilla said there was enough for another dinner and told me to set it on the pantry shelf and cover it. I meant to cover it just as much as could be, Diana, but I forgot all about it. I thought of it next morning and ran to the pantry. Diana, fancy if you can my extreme horror at finding a mouse drowned in that pudding sauce! I lifted the mouse out with a spoon and threw it out in the yard. Marilla was out milking and I fully intended to ask her when she came in if I’d give the sauce to the pigs; but when she did come in I was imagining that I was a frost fairy going through the woods turning the trees red and yellow, so I never thought about the pudding sauce again. Adapted from L. M. Montgomery, Anne of Green Gables What does this passage suggest about Anne? Choices: [(a) She does not take life very seriously and wishes that others would relax., (b) She has good intentions, but she can be absentminded sometimes., (c) She is self-confident and does not worry about what other people think.] Answer index: 1
Subject: Language Arts Skill: analyze-passages-from-anne-frank-the-diary-of-a-young-girl-part-2 Description: Review the passage. Think of it, having to sit in such terror for a day and two nights! We thought of nothing, but simply sat there in pitch darkness—in her fear, Mrs. van D. had switched off the lamp. We whispered, and every time we heard a creak, someone said, "Shh, shh." From Anne Frank, The Diary of a Young Girl: The Definitive Edition. Copyright 1991 by The Anne Frank Fonds. Trans. Susan Massotty What is the meaning of **engrossed** as it is used in the passage? Choices: [(a) preoccupied or captivated, (b) puzzled or confused, (c) absentminded or forgetful] Answer index: 0